Calculate the focusing power of the lens in the fish eye. Assume that the lens is spherical
with a diameter of 2 mm. The index of refraction for water is 1.33, the lens index of refraction 1.4,
and vitreous humor and aqueous humor have similar refrection index 1.34.
1
Expert's answer
2020-09-21T08:32:17-0400

"D = 2 \;mm"

"R = \frac{D}{2} = 1 \;mm = 0.1 \;cm"

"M_{lens} = 1.4"

"M_{ah} = 1.34"

"R_1 = +R"

"R_2=-R"

"\frac{1}{f} = (\frac{M_{lens}}{M_{ah}} -1)(\frac{1}{R_1} – \frac{1}{R_2})"

"\frac{1}{f} = (\frac{1.4}{1.34} – 1)(\frac{2}{R})"

"P = \frac{1}{f} = 0.0447\times20=0.894 \;Dioptre"

Answer: 0.894 Dioptre